### v1

# LaST0: Latent Spatio-Temporal Chain-of-Thought for Robotic Vision–Language–Action Model

Abstract
[…]
I Introduction
Recently, robot imitation learning has emerged as a dominant paradigm for acquiring manipulation policies from expert demonstrations [25, 16, 72]. By inheriting the semantic understanding and common-sense reasoning capabilities of Vision–Language Models (VLMs) [1, 40, 75, 21], Vision–Language–Action (VLA) models integrate rich pretrained knowledge with the low-level control capabilities of robotic policies [6, 5, 50, 35]. This integration endows robotic agents with a unified framework for interpreting human instructions and executing corresponding manipulation primitives in dynamic environments.
Rather than simply mapping observations to actions, recent advances in VLA models have been inspired by the Chain-of-Thought (CoT) reasoning paradigm in general VLMs [30, 89]. In this line of work, some approaches enhance manipulation stability and interpretability by explicitly generating linguistic reasoning traces or affordance representations [85, 46]. In parallel, other studies seek to capture environmental dynamics by predicting future states, such as subgoal images or camera-view depth maps [71, 90, 77, 92, 52]. Despite their demonstrated benefits, explicit CoT VLA methods remain constrained by two fundamental challenges in robotic manipulation. On the one hand, explicit reasoning typically incurs non-negligible inference latency. The autoregressive generation paradigm introduces inevitable computational overhead [69, 76], limiting the VLA model’s ability to achieve real-time responsiveness. This latency further restricts VLA models from reasoning effectively along the temporal dimension, thereby undermining the temporal consistency required for closed-loop manipulation. On the other hand, explicit reasoning is often confined to the linguistic space, imposing a representational bottleneck that struggles to faithfully capture ineffable physical attributes. In contrast, robotic agents must reason about and interact with the physical world, which is essential for robust manipulation in a dynamic environment.
In this paper, we propose LaST0, a dual-system VLA model that enables efficient reason-before-act behavior through a Latent Spatio-Temporal Chain-of-Thought (CoT). As shown in Fig. LABEL:fig:intro, unlike prior explicit CoT-based VLA methods, LaST0 performs reasoning in a compact latent space, enabling the capture of fine-grained physical and robotic dynamics that are difficult to verbalize, while supporting temporally coherent modeling for closed-loop manipulation. Specifically, we introduce a token-efficient latent CoT space that autoregressively predicts future latent tokens of 2D images, 3D point clouds, and proprioceptive states. As a result, the VLA model can implicitly model the semantic and geometric structure of physical dynamics, while forming an internal representation of the robot state, thereby capturing the relationship between the robot and its interactive environment. Meanwhile, the latent CoT space is extended across future keyframes, enabling temporally consistent causal reasoning, which improves action coherence in closed-loop robotic manipulation. Although the proposed latent CoT is compact and encodes richer physical information, incorporating it into action generation still introduces additional inference overhead. Therefore, leveraging temporally extended latent conditions, we further propose a dual-system architecture implemented via a Mixture-of-Transformers (MoT) design, as shown in Fig. LABEL:fig:intro. Specifically, two experts are integrated within a single VLA model: a slow reasoning expert, which performs low-frequency latent inference to capture spatio-temporal dependencies, and a fast acting expert, which generates actions conditioned on high-frequency observations and periodically updated latent representations. Through shared self-attention mechanisms, LaST0 enables long-context interaction between the latent CoT space and the action space, thereby effectively coordinating deliberative reasoning with responsive control.
For the training procedure, both the latent reasoning expert and the action expert are initialized from the same pretrained multimodal foundation model (i.e., Janus-Pro [12]). We then perform large-scale pretraining of LaST0 on diverse robotic manipulation datasets [62, 41, 82], ensuring a shared representation space and seamless interaction between the two experts within a unified VLA model. During downstream training, we first optimize the slow reasoning expert via supervised fine-tuning using ground-truth latent representations extracted from auxiliary encoders. This latent-space teacher forcing establishes a reliable latent scaffold for subsequent action learning. Afterward, the reasoning expert is frozen, and the action expert is trained under heterogeneous fast–slow operating ratios, enabling the VLA model to adaptively select different frequencies during deployment. For evaluation, we systematically assess LaST0 on 10 simulated tasks and six complex real-world tasks, including single-arm and dual-arm settings. Across these benchmarks, LaST0 outperforms prior state-of-the-art (SOTA) VLA methods by 8% and 13%, respectively, while achieving a 14× speedup over previous explicit CoT VLA approaches. In addition, we validate the closed-loop manipulation capability of LaST0 on long-horizon real-world tasks under dynamic conditions. Notably, the robot is able to repeatedly and continuously scoop eggs from a pan while adapting to dynamic environmental changes, demonstrating robust spatio-temporal understanding. Our contributions are summarized as follows:
  • We propose LaST0, a unified VLA model that enables efficient reason-before-act behavior through a Latent Spatio-Temporal CoT, performing reasoning in a compact latent space to capture fine-grained physical and robotic dynamics that are difficult to verbalize.
  • We design a spatio-temporal latent CoT space for manipulation, which autoregressively models future semantic, geometric, and proprioceptive information, allowing LaST0 to reason about physical dynamics in a temporally coherent manner.
  • We introduce a dual system VLA architecture, implemented via MoT scheme, that coordinates low-frequency latent reasoning with high-frequency action generation, enabling real-time robotic manipulation.
II Related Work
II-A Vision-Language-Action Model
Vision-Language-Action (VLA) models have rapidly advanced as an effective paradigm for instruction-conditioned robotic manipulation, primarily driven by scaling robot demonstration data and adapting pretrained VLMs for robotic control [7, 6, 3, 48, 42]. To improve expressivity for continuous actions, recent VLA research has increasingly employed continuous generative policy heads. Diffusion-based VLA [51, 79, 81, 50] models complex action distributions through iterative denoising, while flow-matching formulations [5, 35, 22, 4, 80] offer an alternative that can improve sampling efficiency and stability. These directions aim to leverage the generalization of VLMs to specifically enhance the controllability, precision, and fidelity of continuous action generation. Beyond imitation learning, more work explores post-training VLA models via Reinforcement Learning (RL) to enhance robustness. Offline methods leverage preference-based alignment [91] to steer policies using static data without requiring simulation. Online RL methods [70, 15, 14, 44, 54] refine policies through active environment interaction. However, scaling RL-based strategies for open-world manipulation remains constrained by challenges in sample efficiency and training stability. Meanwhile, recent research equips VLA models with “reason-before-act” components to retain accurate representations of the dynamic environment and improve action quality.  [35, 49, 86] adopts textual chain-of-thought (CoT) generation for future task planning and action generation. Subsequent work [78, 92, 10, 26, 8, 90] extends generative text planning to future image prediction. Furthermore, in order to model fine-grained future states, [52, 9, 28] introduces predictions of future multi-modal information. However, explicit reasoning typically incurs non-negligible inference latency, thereby limiting the VLA model’s ability to achieve real-time responsiveness. Unlike explicit CoT-based VLA methods, LaST0 employs a latent spatio-temporal CoT strategy to efficiently capture fine-grained physical and robotic dynamics within a compact latent space.
II-B Latent Chain-of-Thought
Recent work of VLM in the general domain [13, 23, 31, 84, 76] has explored latent CoT reasoning to address the limitations of explicit CoT on ineffable visual-spatial matching and high-cost text or image generation. These methods perform multi-step inference directly in continuous latent spaces, allowing intermediate reasoning to be compact, implicit, and tightly integrated with downstream prediction. Beyond general-purpose VLMs, similar approaches have been adopted in the embodied intelligence domain, where textual CoT is particularly ill-suited for representing low-level signal generation. LCDrive [69] replaces language-based explanations with action-aligned latent rollouts, allowing the model to simulate future outcomes and evaluate alternatives directly in the action space. Meanwhile, Thinkact [34] compresses intermediate motion plans into compact representations to guide action generation. Prior latent CoT frameworks mainly operate on textual reasoning or abstract 2D representations for general tasks. In contrast, LaST0 is tailored for robotic manipulation, where the VLA model reasons in a physically grounded latent space that jointly encodes semantic intent, geometric structure, and robot state, thereby capturing the embodied interaction between the robot and its environment.
III Method
Figure 2: Overall framework of LaST0. a) We propose LaST0, a unified VLA model with a dual-system architecture. The model is implemented via a MoT scheme with two experts interacting through shared self-attention. The slow reasoning expert operates at a low frequency, taking visual observations and text as input to construct the LaST CoT by autoregressively predicting spatio-temporal latent tokens, which are stored in the KV cache. The fast acting expert operates at a higher frequency and generates actions via flow matching, conditioned on high-frequency observations and periodically updated latent representations. b) We design a spatio-temporal latent space tailored for the reasoning expert, where pretrained modality-specific encoders extract features from future RGB images, point clouds, and robot states at keyframes. These features serve as ground-truth latent CoT targets for supervising the reasoning expert. c) The training procedure consists of three stages with different parameter update strategies, ensuring a reliable latent and robust action generation.
In Section III-A, we describe the preliminaries of VLA models and contrast explicit and latent Chain-of-Thought (CoT) paradigms in the robotic domain. The proposed LaST0 with a Mixture-of-Transformers (MoT) architecture, featuring dual experts coupled via a shared attention mechanism, is detailed in Section III-B. Section III-C elucidates the construction of our spatio-temporal latent space, explaining the encoding and compression of future physical knowledge. We then describe the fast–slow system design and its asynchronous inference frequencies in Section III-D. Finally, Section III-E outlines the complete training paradigm, from large-scale pretraining to supervised fine-tuning.
III-A Preliminaries
VLA Problem Formulation. We formulate the robot manipulation task as a probabilistic sequence decision-making problem [42]. At each timestep $t$, the policy receives a natural language instruction $l_{t}$ and visual observations $I_{t}\in\mathbb{R}^{H\times W\times 3}$ that capture the current environment. The objective of the VLA model $\pi_{\theta}$ is to generate an optimal action sequence $\mathbf{a}_{t:t+H}$ conditioned on the instruction $l_{t}$. Consistent with prior works[50, 5], we define the action space within the Special Euclidean group $SE(3)$. For single-arm configurations (e.g., Franka research 3), we employ a 7-DoF end-effector pose control mechanism, formulated as $\mathbf{a}_{t}\in\mathbb{R}^{7}$. Specifically, this control vector consists of 3-DoF for relative positional offsets ($\Delta x,\Delta y,\Delta z]\in\mathbb{R}^{3}$), 3-DoF for rotation (represented as Euler angles [roll, pitch, yaw] $\in\mathbb{R}^{3}$), and 1-DoF for the gripper state (open/closed, $g\in\mathbb{R}^{1}$). For dual-arm configurations, to validate our model’s scalability, we extend this to a 14-DoF representation by concatenating the control signals.
Robotic CoT Reasoning. To bridge the gap between high-level visual observations and low-level control, recent VLA methods introduce an intermediate CoT variable $\mathcal{Z}$. This formulation decomposes the policy distribution into a reasoning stage (predicting $\mathcal{Z}$) and an execution stage (predicting $\mathbf{a}$ conditioned on $\mathcal{Z}$):
$p(\mathbf{a},\mathcal{Z}\mid I_{t},l)=p(\mathbf{a}\mid\mathcal{Z},I_{t},l)\cdot p(\mathcal{Z}\mid I_{t},l)$
Existing methods typically adopt an explicit CoT paradigm, where $\mathcal{Z}$ consists of discrete natural language tokens [45] or image tokens [92, 88]. While interpretable, these approaches confine reasoning to the linguistic space of LLMs, which struggles to represent ineffable physical attributes and incurs inference latency due to explicit decoding. To overcome these limitations, we propose a new latent CoT paradigm for the robotic domain. Unlike explicit CoT, we define $\mathcal{Z}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{k}\}$ as a sequence of continuous embeddings in a high-dimensional latent space. In our framework, the latent variable $\mathcal{Z}$ is trained to autoregressively predict future dynamics, including latent representations of 2D images, 3D point clouds, and robot proprioceptive states, thereby modeling the physical world in a compact space.
III-B LaST0 Architecture
Overview. As illustrated in Fig. 2 a), we elaborate on the architectural of LaST0. Our model is initialized on the Janus-Pro[12], utilizing DeepSeek-LLM 1B as the backbone. To bridge the gap between reasoning and control, we transform this standard decoder-only transformer into a unified MoT dual-system architecture, constructing the ”reason-before-act” paradigm. This design enables the model to effectively decouple the generation of slow, high-level latent reasoning from fast, low-level action execution, while maintaining seamless information flow through a shared attention mechanism.
Vision Encoder. For each input RGB observation $I_{t}\in\mathbb{R}^{H\times W\times 3}$ ($H=W=224$), we employ SigLIP-Large[87] to extract semantic features. Following standard practices, the image is divided into non-overlapping patches of size $14\times 14$, which are then flattened and linearly projected. This process yields a compact feature sequence $f_{\text{img}}\in\mathbb{R}^{B\times N_{\text{img}}\times d_{v}}$, where $B$ denotes the batch size, $N_{\text{img}}$ represents the sequence length, and $d_{v}$ is the embedding dimension. In our framework, these encoded features $f_{\text{img}}$ serve a dual purpose: the current frame acts as real-time contextual input to the MoT experts, while future frames provide ground-truth target embeddings for the visual component of the latent CoT.
Point Cloud Encoder (training only). To equip the model with fine-grained geometric perception and robust 3D spatial reasoning, we integrate a preyrained point cloud encoder that explicitly captures object geometry and spatial knowlegde. We adopt Uni3D [94], a scalable 3D foundation model, which has been pre-trained on massive 3D datasets to align point cloud features with image and text embeddings. Note that, unlike the vision encoder, the point cloud encoder is not used during inference. Instead, Uni3D is solely employed to encode ground-truth point clouds into compact 3D feature representations within the latent CoT space.
Mixture-of-Transformers LLM Backbone. We adopt DeepSeek-LLM 1B as our foundation, repurposing its 24-layer decoder-only transformer architecture into a unified dual-system policy. Inspired by [20], to efficiently support both high-level reasoning and low-level execution, we transform the standard backbone into a MoT architecture. Unlike conventional transformers that apply a homogeneous set of weights to all tokens, our MoT design introduces task-specific parameter sets for all non-embedding components, including Feed-Forward Networks (FFN), attention projections ($W_{Q},W_{K},W_{V},W_{O}$), and Layer Normalizations, while maintaining a shared global self-attention context. This modification essentially yields two specialized experts residing within the same $d=2048$ dimensional latent space, while a slow reasoning expert responsible for autoregressively synthesizing the Latent CoT embeddings $\mathcal{Z}$ from language and slow-stream visual inputs, and a fast acting expert dedicated to generating high-frequency actions $\mathbf{a}_{t}$. The design of different operating frequencies for the two experts is described in Section III-D.
MLP Components. To further clarify the LaST0 architecture, we describe the remaining auxiliary components, all of which are implemented as Multi-Layer Perceptrons (MLPs). First, a 3D Projector aligns point cloud features to the same dimension, serving as supervise targets for the Latent CoT. Regarding the action generation mechanism, given that our fast acting expert adopts a Flow Matching policy [5], specific modules are incorporated to handle the continuous generative dynamics. For the action expert input, we use a timestep MLP to encode the continuous time coordinate $t\in[0,1]$, initialized with sinusoidal embeddings, and a noised-action MLP to project the perturbed action state. For the action expert output, a projector MLP is used to transform the predicted flow velocity field, enabling the model to generate precise actions.
III-C Latent Spatio-Temporal Chain-of-Thought
To capture fine-grained physical and robotic dynamics that are difficult to verbalize, while enabling efficient temporal modeling for manipulation, we construct a Latent Spatio-Temporal Chain-of-Thought (LaST CoT).
Latent Embedding Construction. To model temporal environmental dynamics, our latent representation encodes multimodal future states over a horizon $H$. As shown in Fig. 2 b), for each future timestep $k\in\{1,\dots,H\}$, we extract features from three complementary modalities to form a holistic physical representation. Future RGB frames $I_{t+k}$ are encoded into visual latents $z^{v}_{k}$ using the frozen SigLIP-Large encoder; simultaneously, future point clouds $P_{t+k}$ are processed by the Uni3D encoder to yield geometric latents $z^{p}_{k}$ capturing 3D spatial occupancy, while future robot states $\mathbf{s}_{t+k}$ are transformed into proprioceptive latents $z^{s}_{k}$ via an action tokenizer. Together, these representations enable the VLA model to implicitly model the semantic and geometric structure of physical dynamics, while maintaining an internal estimate of the robot’s state, thereby capturing the interaction between the robot and its environment. To ensure high inference efficiency, we apply average pooling to compress the feature maps of each modality into a single representative token. This results in a compact set of embeddings $\{z^{v}_{k},z^{p}_{k},z^{s}_{k}\}$ for each step. We then organize these tokens in an interleaved, chronological order to preserve causal physical dependencies:
$\mathcal{Z}_{\text{GT}}=[\mathbf{z}^{v}_{1},\mathbf{z}^{p}_{1},\mathbf{z}^{s}_{1},\mathbf{z}^{v}_{2},\mathbf{z}^{p}_{2},\mathbf{z}^{s}_{2},\dots,\mathbf{z}^{v}_{H},\mathbf{z}^{p}_{H},\mathbf{z}^{s}_{H}].$ (1)
By fusing visual information, 3D geometric structure, and robot proprioception, the model develops a holistic understanding of both the physical world and the robot itself. The interleaved multimodal structure further encourages the model to learn the coupled dynamics across different modalities over time. Notably, the temporal granularity can be flexibly adjusted: we either adopt keyframe extraction as in [67] or use densely sampled frames, depending on task requirements. Moreover, by compressing high-dimensional sensory inputs into a latent sequence of length $3\times H$, we avoid the prohibitive cost of decoding pixel-level images or long textual sequences. As discussed in Section III-D, we introduce an asynchronous frequency design for the two experts, which further accelerates action generation.
Sequence Structure. To better organize LaST CoT reasoning and action generation, we introduce three special tokens: <latent_start>, <latent_end>, and a placeholder token <latent_pad>. The reasoning segment is structurally defined as a sequence bounded by the start and end tokens, with the intermediate positions reserved for the latent embeddings. During Training, we replace the intermediate <latent_pad> tokens with the ground-truth latent sequence $\mathcal{Z}_{\text{GT}}$. This allows the model to learn the transition dynamics via standard teacher forcing. While during Inference, the model is initialized with <latent_start> followed by a sequence of <latent_pad> tokens. The slow reason expert then autoregressively generates continuous embeddings, sequentially filling the positions of the <latent_pad> placeholders until the pre-defined horizon is filled. The horizon length can be adaptively adjusted, and its effect is further analyzed through ablation studies.
Latent Supervision Strategy. We train the slow reasoning expert to predict latent CoT embeddings in an autoregressive manner, using continuous latent regression rather than discrete token likelihoods [84, 76]. Specifically, the slow expert is trained to predict a sequence of latent reasoning states $\hat{\mathcal{Z}}$ in a next-step prediction manner, conditioned on preceding observations and context. Unlike conventional CoT supervision based on discrete token prediction, our latent targets consist of continuous, high-dimensional embeddings that encode future physical world states. To align the predicted latent semantics with the ground-truth representations, we employ cosine similarity as the supervision objective. The loss is defined as:
$\mathcal{L}_{\text{latent}}=\sum_{t=1}\left(1-\frac{\hat{\mathbf{z}}_{t}\cdot\mathbf{z}_{t}^{\text{GT}}}{\lVert\hat{\mathbf{z}}_{t}\rVert\,\lVert\mathbf{z}_{t}^{\text{GT}}\rVert}\right).$ (2)
By maximizing directional alignment in the latent space, this objective encourages the model to anticipate future physical dynamics in a structured and compact manner. The slow expert learns a physically grounded latent reasoning trajectory that serves as a stable scaffold for the action generation.
III-D Dual-System Coordination
Figure 3: The slow reasoning expert performs low-frequency latent CoT reasoning to capture long-horizon spatio-temporal dependencies, while the fast acting expert generates actions conditioned on high-frequency observations and periodically updated latent knowledge. Benefiting from training with a mixture of fast–slow operating ratios, the model can flexibly adjust the reasoning–action frequency ratio at test time.
Asynchronous Frequency Coordination. To coordinate LaST CoT reasoning with high-frequency robotic control, we introduce an asynchronous frequency mechanism between the slow reasoning expert and the fast acting expert. As shown in Fig. 3, we decouple their operating frequencies by a fixed update ratio $\kappa$ (e.g., $\kappa\in{2,4,8}$). The slow expert is activated only at sparse keyframes ($t\bmod\kappa=0$), where it performs autoregressive latent CoT reasoning. In contrast, the fast expert runs at the native control frequency and remains active at every timestep. Between consecutive keyframes ($t\bmod\kappa\neq 0$), the slow reasoning expert is held dormant, while the fast expert generates actions by conditioning on the most recent latent reasoning output. This strategy ensures that action generation is continuously guided by stable, temporally consistent latent conditions, while minimizing redundant computations for LaST CoT reasoning.
For the inputs to the two experts, the slow reasoning expert receives the natural language instruction $l$ and the low-frequency observation $I_{\text{slow}}$, constructing the Latent CoT that encapsulates future physical dynamics. Conversely, the fast acting expert is optimized for rapid closed-loop feedback and receives only the high-frequency observation $I_{\text{fast}}$. Crucially, since our MoT architecture maintains a unified token sequence, the fast expert can efficiently attend to both the linguistic goal and the Latent CoT tokens via the shared attention mechanism.
Inference via KV Cache. During inference, the Key-Value states computed by the slow expert (encapsulating the Latent CoT) are cached in memory. During intermediate fast-control steps, the acting expert only encodes the current observation and attends to the frozen latent CoT cache, effectively retrieving the CoT tokens with $O(1)$, without re-invoking the slow reasoning process. This approach eliminates repetitive decoding, allowing our model to achieve an inference speed of 15.4 Hz on a single RTX 4090 GPU when operating with a 1:4 fast–slow frequency ratio.
III-E Training Recipe
Large-Scale Robotic Pretraining. Both the latent reasoning expert and the action expert are initialized from the Janus-Pro 1B foundation model. As shown in Fig. 2 c), Stage 1 performs large-scale pretraining on a diverse corpus of over 400K trajectories aggregated from Open-X-Embodiment [62], DROID [41], ROBOMIND [82, 33], and other robotic datasets. Notably, beyond following prior VLA works in applying data quality filtering [42, 52], we additionally ensure that all robot state annotations are accurate and physically consistent. Detailed statistics and descriptions of the pretraining data are provided in Appendix -A. This stage optimizes the model on broad robotic datasets to establish a shared representation space, enabling seamless interaction between reasoning and execution within the unified VLA framework.
Supervised Fine-Tuning (SFT). As shown in Fig. 2, we employ a decoupled 2-stage SFT in downstream adaptation, aiming to prevent the catastrophic forgetting of reasoning capabilities when training action expert. In stage 2, we first optimize the slow reasoning expert while freezing the action expert. By exclusively optimizing the Latent CoT regression loss $\mathcal{L}_{\text{latent}}$, the slow expert aligns its latent representations with domain-specific physical dynamics. Subsequently, in stage 3, we freeze the converged slow expert and load the robotic data pretrained weights for the fast expert. The fast acting expert is subsequently fine-tuned using the standard Flow Matching loss $\mathcal{L}_{\text{flow}}$. While the loss formulation and noise injection scheme follow $\pi_{0}$ [5], our approach differs in that the entire LLM is directly used as the action denoising model, without introducing any additional action heads. Meanwhile, the action expert is trained with randomly mixed fast–slow operating ratios (e.g., 1:1, 1:2, 1:4), which exposes it to latent conditions updated at varying delays. As a result, LaST0 can flexibly accommodate different reasoning-update frequencies at deployment and adaptively choose the fast–slow inference rate. In ablation studies, we find that training with mixed fast–slow operating ratios does not degrade performance; instead, it improves the model’s robustness during inference.
IV Experiment
In Section IV-A, we evaluate the manipulation performance and inference efficiency of LaST0 against prior methods in simulated environments. The contribution of each component is analyzed through ablation studies in Section IV-B. Section IV-C presents both quantitative and qualitative results of LaST0 on real-world manipulation tasks, including single-arm and dual-arm control under diverse robot configurations.
TABLE I: Comparison of LaST0 and baselines on RLBench. All methods are trained in the multi-task setting [67], and we report mean success rates (S.R.). Inference speed is evaluated on an NVIDIA 4090 GPU.
Models Close Close Toilet Sweep Close Phone Umbrella Frame Wine at Water Mean S.R. $\uparrow$ Infer.
box laptop lid seat down to dustpan fridge on base out off hanger rack plants & Var $\downarrow$ speed $\downarrow$
OpenVLA [42] 0.60 0.35 0.75 0.55 0.85 0.20 0.30 0.15 0.20 0.05 0.40 $\pm$0.02 6.3 Hz
SpatialVLA [64] 0.80 0.70 0.85 0.20 0.80 0.15 0.25 0.40 0.15 0.30 0.46 $\pm$0.03 7.9 Hz
CogACT [47] 0.90 0.80 0.95 0.50 0.85 0.50 0.55 0.45 0.30 0.25 0.61 $\pm 0.04$ 9.8 Hz
CoT-VLA [92] 0.95 0.75 1.00 0.80 0.65 0.50 0.40 0.50 0.55 0.50 0.66 $\pm 0.03$ 1.1 Hz
$\pi_{0.5}$ [35] 0.90 0.95 0.85 0.75 1.00 0.05 0.10 0.80 0.75 0.35 0.65 $\pm$0.04 13.8 Hz
HybridVLA [50] 0.85 0.95 1.00 0.90 1.00 0.50 0.50 0.70 0.50 0.50 0.74 $\pm$0.04 6.1 Hz
LaST0 0.95 0.95 1.00 0.80 0.85 0.75 0.75 0.70 0.85 0.65 0.82 $\pm$0.03 15.4 Hz
Figure 4: Ablation study on key design choices of LaST0. We analyze (a) the importance of different latent modalities, (b) the number of tokens allocated per latent modality, (c) the temporal coverage in latent reasoning, and (d) the collaboration frequency between reasoning and action experts. Results are reported as average success rates across 10 RLBench tasks, demonstrating the contribution of each component to overall performance.
IV-A Simulation Experiment
Data collection. We evaluate LaST0 on a diverse set of 10 manipulation tasks from the RLBench [36] benchmark, conducted in the CoppeliaSim simulation environment. The evaluated tasks include 1) Close box, 2) Close Laptop, 3) Toilet seat down, 4) Sweep to dustpan, 5) Close fridge, 6) Phone on base, 7) Take umbrella out, 8) Frame off hanger, 9) Wine at rack, and 10) Water plants. All tasks are executed using a Franka Panda robotic arm with a single front-view observation. Demonstration data are collected by following pre-defined waypoints with motion plans generated via the Open Motion Planning Library [68]. Following the frame-sampling protocol adopted in prior work [67, 27, 38], we construct a training dataset comprising 100 demonstration trajectories per task.
Training and evaluation protocol. We benchmark LaST0 against six representative state-of-the-art VLA models: OpenVLA [42], $\pi_{0.5}$ [35], CogACT [47], SpatialVLA [64], CoT-VLA [92] and HybridVLA [50]. For all baselines, we initialize from the officially released pretrained checkpoints and follow the full fine-tuning configurations recommended by the paper. For CoT-VLA, we reimplement the method on top of Janus-Pro [12], the same foundation model used in our approach, and reproduce its explicit CoT reasoning to ensure a fair comparison. For LaST0, each observation consists of a single RGB image resized to $384\times 384$, a point cloud uniformly subsampled to 1024 points, a language instruction obtained directly from the simulator, and the robot proprioceptive state synchronized with the predicted action. Note that point cloud data are used exclusively during training to construct latent ground truth and are not needed during inference. The model is trained for 40 epochs in Stage 2 and 300 epochs in Stage 3, using the AdamW optimizer [53] across 8 NVIDIA A800 GPUs. Evaluation follows the protocol of prior work [47, 27]. For each task, we perform 20 rollout trials using the final checkpoint, repeat the evaluation three times with different random seeds, and report the mean success rate together with its variance.
Quantitative analysis. As summarized in Table I, LaST0 achieves a mean success rate of 82% across 10 RLBench manipulation tasks, substantially outperforming all prior baselines. In particular, LaST0 surpasses the strongest existing methods, HybridVLA-7B (74%), $\pi_{0.5}$-3B (65%), and CogACT-7B (61%), by margins of 8%, 17%, and 21%, respectively. Beyond the overall average, LaST0 attains the highest success rate on 7 out of 10 tasks, indicating consistent performance gains across diverse manipulation skills, while also showing relatively low variance ($\pm$0.03), reflecting stable execution across tasks. These improvements primarily stem from the latent reasoning representations produced by the reasoning expert, which condition the action expert with compact latent states encoding future visual dynamics, 3D spatial structure, and robot proprioceptive information, enabling more stable and temporally coherent action generation. In terms of execution efficiency, LaST0 operates at an inference speed of 15.4 Hz with an action chunk size of one, which is significantly faster than explicit CoT methods (CoT-VLA: 1.1 Hz) and remains competitive with $\pi_{0.5}$ (13.8 Hz). This demonstrates that LaST0 achieves superior task performance without sacrificing real-time control efficiency. Moreover, as shown in Fig. 5, we compare the attention heatmaps of LaST0 against variants without CoT and with explicit CoT (CoT-VLA). While the no-CoT variant and the explicit CoT method fail to aggregate features on the manipulated objects and ignore the relationship between the robot and the manipulated objects, LaST0 exhibits a highly concentrated attention pattern, highlighting its superior understanding of spatio-temporal information in downstream tasks.
TABLE II: Comparison of LaST0 and baselines in real-world manipulation tasks. We report success rates (S.R.) for standard single-arm and dual-arm tasks (left), and a long-horizon task evaluated over three consecutive steps (right).
Models Wipe Press Place dish Place egg Scoop popcorn Open pot Mean Place egg on bread (Long-horizon)
blackboard stamp on rack on bread into bowl pick corn S.R. $\uparrow$ Step 1  $\rightarrow$ Step 2    $\rightarrow$ Step 3
SpatialVLA [64] 0.60 0.67 0.30 0.20 0.27 0.40 0.39 0.20 0.07 0.00
$\pi_{0.5}$ [35] 0.60 0.73 0.60 0.47 0.53 0.60 0.59 0.47 0.20 0.07
CoT-VLA [92] 0.53 0.60 0.66 0.53 0.33 0.53 0.53 0.33 0.13 0.07
LaST0 0.73 0.93 0.80 0.66 0.66 0.53 0.72 0.66 0.47 0.33
IV-B Ablation Study
To validate the key design choices of LaST0, we conduct comprehensive ablation experiments on 10 RLBench tasks, using the same training and evaluation settings as in the simulation experiments. The results are summarized as follows.
Importance of latent CoT modalities. As shown in Fig. 4 a), we analyze the contribution of each latent modality by selectively ablating individual latent components while keeping the overall architecture unchanged. When using only the image, point cloud, or robot state latent, the model achieves success rates of 74%, 76%, and 75%, respectively, indicating that each modality-specific latent provides a strong basis for action generation. The combination of multiple modality latents continues to provide additional performance improvements, even when the manipulation accuracy is already high. These results validate the importance of modeling comprehensive physical dynamics in the latent space, and further demonstrate that enabling the model to autonomously reason about the relationship between the robot and its interactive environment is effective for robotic manipulation.
Figure 5: Attention heatmap visualizations from the last layer for three VLA models: (a) LaST0 without CoT reasoning, (b) the explicit CoT in CoT-VLA, and (c) LaST0 with LaST CoT.
Number of tokens per latent modality. As shown in Fig. 4 b), we study how the token budget allocated to each latent modality affects performance by varying the number of latent tokens while keeping other components fixed. When no latent token is used, performance drops substantially to 68%, indicating that the absence of latent representations severely limits the model’s reasoning capacity. Introducing a single token for each modality leads to a sharp improvement (up to 82%), demonstrating that even a minimal latent representation is sufficient to establish an effective latent decision state.
Further increasing the number of tokens yields no significant accuracy improvement, suggesting that high-level latent tokens can compactly encode physical information sufficient for effective reasoning.
Temporal coverage in latent reasoning. As shown in Fig. 4 c), we investigate the effect of the temporal horizon used in latent reasoning by varying the number of future time steps encoded into the latent state. Performance improves consistently as the latent temporal coverage increases (from 68% to 82% when extending from 0 to 4 steps), indicating that incorporating longer temporal dependencies enables more informed latent decision states. At the same time, expanding the temporal horizon introduces higher inference latency, revealing a trade-off between decision quality and computational efficiency in reasoning expert. Due to our fast–slow system design, extending the temporal horizon of the latent space does not significantly affect action generation speed.
Collaboration frequency between reasoning and action experts. As shown in Fig. 4 d), varying the collaboration frequency has a clear impact on task success. Fixed ratios such as 1:1, 1:2, and 1:4 achieve comparable performance (75-79%), while overly infrequent collaboration (1:8) leads to a relative drop to 74%. Finally, the mixed strategy trained by combining data from all four collaboration ratios achieves the best performance (82%), using a 1:4 ratio during testing. These results indicate that our frequency joint training scheme enables more robust coordination between the reasoning and action experts across tasks.
Figure 6: Real-world Experiment Setup. We show the details about robot-arm setup and assets of manipulation tasks.
Figure 7: Visualization of task execution processes by real-world single-arm and dual-arm robots (from left to right).
IV-C Real-World Experiment
Data collection. We evaluated our method on a set of real-world manipulation tasks using both single-arm and dual-arm Franka robot setups. As shown in Fig. 6, the single-arm platform consists of a Franka robot equipped with two Intel RealSense D455 cameras, providing third-person and wrist-view RGB-D observations. The dual-arm setup combines two Franka robots and three D455 cameras, including one third-person view and two wrist-mounted views. In the single-arm setting, we collect demonstrations for three tasks: 1) wiping a whiteboard with an eraser, 2) placing a dish on a rack, and 3) placing an egg on bread using a spatula, which is a complex task requiring the robot to precise scoop the egg and place it back onto the bread. In addition, we further evaluate a long-horizon setting on placing egg task, where the robot consecutively completes the full task three times while the positions of the manipulated objects continuously change during execution. For the dual-arm setting, two collaborative tasks were considered: 1) scooping popcorn into a bowl and 2) opening a pot lid followed by picking corn from the pot. All demonstrations were collected using the Gello platform [83], with 200 high-quality demonstrations per task. More details are provided in Appendix -B and Appendix -C.
Training and evaluation details. We train the LaST0 policy following the same protocol as in simulation, with the only difference being the use of multi-view visual inputs in real-world settings: two camera views for the single-arm setup and three views for the dual-arm setup. We compare our method against three strong baselines: $\pi_{0.5}$ [5], a SOTA 2D VLA model; SpatialVLA [64], a SOTA 3D VLA model; and CoT-VLA [92], an explicit CoT-based VLA model that enhances action generation by predicting future visual observations. For fair comparison, all models use the same number of camera viewpoints, and each task is evaluated with 15 rollouts under identical test conditions.
Quantitative and qualitative analysis. As shown in Table II, LaST0 achieves the best overall performance on real-world manipulation tasks, with a mean success rate of 72%, substantially outperforming SpatialVLA (39%), $\pi_{0.5}$ (59%), and CoT-VLA (53%). LaST0 consistently delivers strong gains across a diverse set of tasks, particularly those requiring precise spatial reasoning and temporally coherent control. For instance, on Place dish on rack, LaST0 reaches a success rate of 80%, compared to 66% for CoT-VLA and 60% for $\pi_{0.5}$. On Press stamp, LaST0 further achieves a success rate of 93%, significantly outperforming $\pi_{0.5}$ (73%) and CoT-VLA (60%). These results highlight the advantage of LaST0’s latent spatio-temporal reasoning in modeling physical dynamics and robot-environment interactions in real-world settings.
[…]
Fig. 7 illustrates representative task executions across single-arm and dual-arm settings. We observe that LaST0 produces smooth and continuous motions, particularly in precise actions such as surface wiping, spatula-mediated placement, and bimanual scooping. This behavior is enabled by the proposed LaST CoT and dual-system design, where a low-frequency reasoning expert provides temporally coherent latent guidance, while the action expert operates at a higher control frequency, allowing rapid and fine-grained response to environment dynamics. We also verify the model’s closed-loop capability: for example, when new marks are continuously drawn on the blackboard during manipulation, the robotic arm can persistently perform the erasing action.
V Conclusion and Future Work
In this work, we introduced LaST0, a dual-system VLA model that enables efficient reason-before-act behavior for robotic manipulation through a Latent Spatio-Temporal Chain-of-Thought (LaST CoT). By shifting reasoning from explicit traces to a compact latent space, LaST0 overcomes the latency and representational bottlenecks inherent in prior explicit CoT VLA approaches, while preserving the ability to model fine-grained physical dynamics essential for closed-loop control. Central to our framework is a token-efficient spatio-temporal latent representation that autoregressively captures future semantic, geometric, and proprioceptive dynamics. This design allows the model to form temporally coherent internal reasoning trajectories, bridging high-level intent understanding with low-level execution. Building upon this LaST CoT mechanism, we further proposed a fast–slow dual-system architecture implemented via a Mixture-of-Transformers, which decouples low-frequency deliberative reasoning from high-frequency action generation. The resulting asynchronous coordination enables real-time responsiveness without sacrificing CoT reasoning capability. We believe LaST0 represents a step toward more scalable and physically grounded reasoning in robotic foundation models.
Future work will further explore richer designs of the latent spatio-temporal reasoning space, enabling more expressive and structured physical abstractions. We also plan to investigate advanced post-training strategies, such as jointly optimizing latent reasoning and action generation via reinforcement learning, as well as scaling LaST0 to more complex long-horizon manipulation tasks with delayed rewards, contact-rich dynamics, and evolving environments.
Acknowledgments